OpenStudy (anonymous):

find binomial expression (2x-y)^4

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

fourth line of pascals triangle is 1 4 6 4 1

OpenStudy (anonymous):

1(2x)^4 + 4(2x)^3(-y) + 6 (2x^2)(-y)^2 + 4(2x)(-y)^3 + (-y)^4 is the first step second step is computation

OpenStudy (anonymous):

stupid typo 1(2x)^4 + 4(2x)^3(-y) + 6 (2x)^2(-y)^2 + 4(2x)(-y)^3 + (-y)^4

OpenStudy (anonymous):

i got nothing.:/ would you please explain 1st step a little?

OpenStudy (anonymous):

1 1 1 1 2 1 1 3 3 1 1 4 6 4 1

OpenStudy (anonymous):

You start with 1 1 1 1 1 1 1 Then you the top two to get the next row: 1 1 1 1 2 1 1 1 since 1 + 1 = 2 1 1 1 1 2 1 1 3 3 1 since 1 + 2 = 3, and 2 + 1 = 3

OpenStudy (amoodarya):